OpenStudy (anonymous):

identify the vertex, focus and directrix of the graph equation. x=1/24y^2

OpenStudy (amistre64):

you need to be able to recall the geometric equation of a parabola

OpenStudy (amistre64):

4ay = x^2 to adapt it to your situation, 4ax = y^2 would seem to fit

OpenStudy (amistre64):

x = 1/24 y^2 24x = y^2 4a = 24, what is a?